OpenStudy (mollykenney):

PLEASE HELP let f(x)= 2x+3 and g(x)= x^2-x 1. (f/g)(2)

OpenStudy (alyssa_xo):

plug 2 into (f/g)(x)

OpenStudy (zale101):

\(\Large\frac{f(2)}{g(2)}=?\)

OpenStudy (alyssa_xo):

\(\large (f/g)(x)=\dfrac{f(x)}{g(x)}=\dfrac{2x+3}{x^2-x} \\ f(2)/g(2)=\dfrac{2(2)+3}{2^2-2}\)

OpenStudy (mollykenney):

thank you so much @alyssa_xo !

OpenStudy (alyssa_xo):

technically thank you zale :p

OpenStudy (mollykenney):

and @Zale101 !

OpenStudy (zale101):

No problem! @alyssa_xo AND @mollykenney